A valószínűségszámítás és matematika statisztika gyakorlaton való részvétel

dátum:  IX. 12. IX. 19. IX. 26. X. 3. X. 10. X. 17. X. 24.

diákok:

Bertók Kornél  +  +  −  −  +
Blaskó János  +  −  +  +
Buka Tamás  +  +  +  +
Fekete Gyula  +  +  −  +  +
Hajnal Gergő  +  +  +  +  +
Harangozó Roland  +  +  +  +  +  +
Horvai Mihály  +  +  +  −
Hudák Levente  +  +  +  +
Kárándi Róbert  +  +  +  +  −
Kenyeres Dávid  +  +  +  +  +  −
Kocsis Dániel László  +  +  +  +  +
Koszczelnik Tamás  +  +  +  −
Kovács Tibor Krisztián  +  +  +  +  −  +
Nagy István  +  −  +  +  +
Ócsai Krisztián  +  +  +  +  +
Orosz Lajos  +  +  −  +  −
Pongor Attila  +  +  +  +  −
Sárkány György  +  +  +  +  +  +
Sepa Zoltán  +  +  −  +  −
Spisák Tamás  +  +  +  +  +  +
Szabó Tamás  +  +  +  +  +  +
Szatmári András  +  +  +  +  +
Takács Péter  +  −  −  −  −
Thold Tibor  +  +  +  +  +  −
Tóth László  +  −  −  −  −  −
Vas Gyula  +  +  +  +  +  +
Zelei Attila  +  +  +  +  +

dátum:  X. 31. XI. 7. XI. 14. XI. 21. XI. 28. XII. 5.

diákok:

Bertók Kornél  +  −  +  +  +  −
Blaskó János  +  +  +  +  +  +
Buka Tamás  +  +  +  +  +  +
Fekete Gyula  −  +  −  −  −  +
Hajnal Gergő  −  +  −  +  +  +
Harangozó Roland  −  +  +  +  +  +
Horvai Mihály  +  −  −  −  −  −
Hudák Levente  +  +  +  +  −  +
Kárándi Róbert  +  −  −  −  −  −
Kenyeres Dávid  +  +  −  +  −  +
Kocsis Dániel László  +  +  +  +  +  +
Koszczelnik Tamás  −  +  +  +  +  +
Kovács Tibor Krisztián  +  +  +  −  +  +
Nagy István  +  +  +  +  +  +
Ócsai Krisztián  +  +  +  +  +  +
Orosz Lajos  +  −  −  +  +  −
Pongor Attila  +  −  −  −  −  −
Sárkány György  −  +  +  +  +  −
Sepa Zoltán  −  +  −  −  +  −
Spisák Tamás  +  +  +  +  +  +
Szabó Tamás  +  +  +  +  +  +
Szatmári András  +  +  +  +  +  +
Thold Tibor  −  +  −  +  +  +
Vas Gyula  +  +  +  +  +  +
Zelei Attila  −  +  −  +  +  +